# Leveraging Photometry for Deconfusion of Directly Imaged Multi-Planet Systems

### v1

@@ 2 Approach
In this work, we investigate the influence of photometry on deconfusion by developing a photometry model and a new likelihood orbit-ranking scheme to pair with the deconfuser. We demonstrate the utility of photometry for differentiating orbit assignments using a subset of highly confused test cases from three inclination groups (low, medium, and high inclination), assuming all planets are visible at all times. We adopt the definition of confusion from P22, where a system is considered to be confused if the deconfuser provides two or more closely ranked orbit assignments per system. When fit by the deconfuser, orbit options may be considered closely ranked if they fall within the acceptable rms error tolerance, which in this work is set to 0.05 AU. After fitting orbits and ranking assignments based on the number of matched detections with the deconfuser, we use an additional ranking scheme to break ties between orbits if they have a more statistically likely brightness result. Figure 2 illustrates the combined flow of the algorithms presented in this work.
@@ 2.1 Photometry Model
The photometry ranking scheme begins with a photometry model at one bandpass, which returns the phase angle and brightness of a planet at any location on its orbit and calculates the brightness measured at the instrument for each detection. To determine the expected contrast of a planet detection, we begin by calculating the planet’s separation from its host star. The deconfuser provides the three-dimensional coordinates (x, y, z) of a simulated planet’s location (see Fig. 2, step 1). These coordinates are used to calculate the planet’s separation from its host star with the basic Euclidean distance formula. Then, we use the coordinates to determine the phase angle of the planet detection with respect to the observer (see Fig. 2, step 2). As described in Section 1, the phase angle is defined as the angle between the host star, exoplanet, and the observer. Given this geometry, the phase angle is calculated using Equation 2,
@@
where $\vec{r}$ is the vector describing the planet’s separation from the host star and $\vec{s}$ is the vector describing the system’s separation from the observer. For the system cases described here (details listed in Table 2), the planetary system locations are set to a distance of ten parsecs, placing them well within the range of distances of the most accessible targets presented in NASA Exoplanet Exploration Program’s Mission Target Star List for HWO (Stapelfeldt & Mamajek, 2023).
@@ 2.1.1 Phase Function
The phase angle calculated in Equation 2 is used to determine the phase function of the planet at each orbital phase. For simplicity, we assume a Lambertian phase function and adopt the following equation (as presented in Brown 2005):
@@
Combining Equations 1 and 3 brings us to an “expected” brightness of an observation. This is the planet-star flux ratio that we would expect to receive at the observing system.
Assuming Lambertian reflectance is a simple approximation that is commonly chosen when modeling and comparing reflected light from planets (e.g., Zugger et al. 2010; Fujii & Kawahara 2012; Cowan et al. 2013; Morgan et al. 2019; De Cock et al. 2022), but the Lambertian assumption neglects the dependence of scattering on albedo, wavelength, and complex atmospheric properties (e.g., Greco & Burrows 2015; Mayorga et al. 2016). It has been demonstrated in the Solar System that real planetary bodies deviate from Lambertian reflectance (e.g., Mallama 2009; Mayorga et al. 2016; Mallama & Hilton 2018; Madden & Kaltenegger 2018). However, for an initial study of how planetary phase variation can inform orbit differentiation, a Lambertian approximates the overall shape of a phase curve well to model phase angle variation. Future work will incorporate phase functions that describe more complex processes, such as the one- or two-term Henyey-Greenstein phase functions, which introduce additional parameters to model forward- and back-scattering processes (Henyey & Greenstein, 1941; Irvine, 1965).
@@ 2.1.2 Noise Model
To consider more realistic observations and to obtain the likelihood of a planet detection, we developed a noise model to include detector parameters in the planet flux from the photometry model (see Fig. 2, step 3). We adopt the instrument parameters defined for the Roman Coronagraph Instrument, as Roman’s Coronagraph has been extensively modeled and tested in preparation for its expected launch in 2026 (Perkins et al., 2024). The instrument parameters for the Roman Coronagraph Instrument are available online through IPAC at Caltech [^3] https://roman.ipac.caltech.edu/sims/Param_db.html, and those we use in our code (from Phase C) are listed in Table 1.
@@ 2.1.3 Detector Model
We convert the planet flux density to the planet photon count rate at the telescope following the method described in Robinson et al. (2016), which was developed with a telescope like Roman in mind and modeled on the work presented in Brown (2005). The detector noise function accepts the calculated planet photon rate for each observation and propagates it through a simplified Electron-Multiplying Charge-Coupled Device (EMCCD) detector model, which adds noise properties to the signal. The model includes shot noise, dark current, read noise, and EMCCD gain. The result is a planet signal with detector noise added for each simulated detection, which amounts to the output electron counts from the detector due to the incoming planet signal. For each orbit option returned by the deconfuser, we calculate both the “observed” noisy photometry and the “expected” noisy photometry of the simulated planet detection. The “observed” photometry simulates the instrument’s observation of a planet. In other words, this is the measured photometry that a user would provide to the algorithm for a real observation. The “expected” photometry is the value one expects to receive from a detected planet based on the phase angle and planet parameters, including detector noise.
@@ 2.2 Photometry Likelihood Ranking
After we have included the detector noise for each observation, we implement a ranking scheme to sort confused orbit options based on the likelihood of the calculated photon rates per system option, given the assigned orbital parameters. We obtain the likelihood of each detection from the probability distribution of the brightness. This value is returned by Figure 2, step 4.
As mentioned in Section 2.1, the detector noise model uses instrument parameters defined for Roman. To generate the probability distribution for each detections’ brightness, the detector noise model first draws samples from a Poisson distribution describing the expected number of photons reaching the detector. The incident photons are multiplied by the detector quantum efficiency before the dark current is added to the signal. We account for the EMCCD gain ($g$) by drawing samples from a Gamma distribution, following the model described in Hirsch et al. (2013) and applied in LOWFSSim [^4] https://github.com/nasa-jpl/lowfssim (the Roman Coronagraph Instrument optical model of low-order wavefront sensing and control; Dube et al. 2021). The shape parameter of the gamma distribution is the number of input photoelectrons ($n_{ie}$), and the scale parameter ($\theta$) is
@@
The contributions to the noise model are a combination of three probability distributions: a Poisson distribution, $P(n_{p})$, describing the incident photons (Equation 5), a Gamma distribution, $\Gamma(n_{ie},\theta)$, describing the EM register of the EMCCD (Equation 6), and a normal distribution, $\mathcal{N}(R_{e^{\mathrm{-}}})$, describing the read noise (Equation 7) to convert the signal to image values. The following equations describe these distributions.
@@
Where $q$ is the detector quantum efficiency, $\Delta t$ is the observation time, and $CIC$ is the detector clock-induced charge (see also Table 1). Combining equations 56, and 7 gives the probability of measuring a certain number of counts for a detection, $p(n_{c})$:
@@
The output of the noise model is the probability density function ($f(y|n_{p})$ in Equation A2) describing the number of photoelectrons read out by the detector, from which we can determine the likelihood. The likelihood of a detection, given the orbital parameters of the orbit fit, is obtained by comparing where the observed brightness falls along this probability density function. We assume each planet detection is an independent event, so the likelihood of measuring all planet detections given the parameters of the matched orbits is the product of each individual detection. The resulting likelihood, $L_{\mathrm{system}}$, is used as a new ranking score for each set of confused orbit options per system. A more detailed mathematical description of this process is given in Appendix A. We note that in some cases, it may be beneficial to examine likelihoods on a per-planet basis to avoid a biased likelihood due to one poorly fit planet. We expand on this in more detail in Section 4.
@@ 2.3 Simulated Systems
This section describes our method for identifying and selecting an example subset of highly confused systems to demonstrate the photometry ranking scheme on and provides the parameters of the simulated systems. We simulate ten test case systems across three inclination groupings, low ($i<45^{\circ}$), medium ($45^{\circ}<i<70^{\circ}$), and high ($i>70^{\circ}$), to demonstrate how exoplanet photometry supports deconfusion of multi-planet systems at all inclinations. The ten simulated systems are each different systems, and only inclination is changing across the three inclination groupings, creating a total of thirty cases for study. Each of the simulated systems contains three planets. We randomly simulated more than two thousand systems to find the ten representative cases used in this work, and chose the first ten from the larger set which were confused across all three inclination groupings. The methods for simulating and identifying these confused systems are described in Section 2.3.1 below.
We restrict this analysis to three-planet systems to demonstrate the photometry ranking scheme based on P22’s findings regarding confusion in systems with three or more planets. The example cases in this work were intentionally chosen because they are clearly challenging cases. That is, each of the simulated systems examined here results in confusion every time they are passed through the initial version of the deconfuser (without photometry), at each of the low, medium, and high inclinations. The purpose of this work is not to present large number statistics, but rather to demonstrate that for these representative known highly confused scenarios, photometry can improve deconfusion. We also demonstrate the photometry ranking scheme on a few Solar System planets in Appendix B.
The simulated systems presented here are not explicitly evaluated for dynamical stability. We simulate systems with orbital separations greater than 0.3 AU, with mutual inclinations and low eccentricities. For planets with masses similar to Earth, separations of 0.3 AU would place planets at separations well outside several times the Hill radii of the other planets in the system (see e.g., Gladman 1993 or Dulz et al. 2020 for discussion of using mutual Hill radii as a basic stability criterion). Dynamical simulations by Smith & Lissauer (2009) indicate that Earth-mass planets in three-planet systems are stable for much longer than one billion years if their initial orbital separation is at least 7.5 times their mutual Hill radii. He et al. (2020) also show that stable multi-planet systems generally have lower median eccentricities and mutual inclinations.
@@ 2.3.1 Identifying Confused Systems
The simulated systems were chosen by first identifying confused systems with inclinations of less than 45. Due to the low probability of confusion for systems with $i<45^{\circ}$, it is more challenging to identify confused low inclination cases. For this reason, we identify test cases for this work starting with low inclination simulations and working our way up. Systems with inclinations less than $\sim$20 are also rarely confused (P22). Thus, there are no systems in our sample set which have an inclination less than 20.
Low inclination systems were simulated with three co-planar planets with random orbital parameters drawn from uniform distributions with semimajor axes between 0.5 and 6.0 AU, eccentricities less than 0.1, inclinations less than 45, and minimum orbital separations between planets of 0.3 AU. Due to the low probability of confusion for systems with low inclination, we simulated two thousand systems to identify highly confused low inclination systems that also resulted in confusion at medium and high inclinations. These systems were simulated using the system generation function (sample_planets.random_planet_elements) contained within the deconfuser framework.
Once a low inclination system was found to be confused with the first version of the deconfuser, that system was passed through the deconfuser again with only the inclination changed. Updated inclination values were randomly sampled from uniform arrays ranging between 45 to <<70 until a confused system was identified in the medium inclination range through empirical trial and error testing. The same method was used for the high inclination grouping, but inclinations were sampled between 70 and 90. For the majority of systems, we tested multiple inclination values in the medium and high groupings to identify a confused system. There were also cases when the orbital parameters for a confused low inclination system did not result in confusion at higher inclinations. In these cases, we removed the low inclination confused system from our sample and replaced it with another low inclination system to build the full sample set. If a system returned confused options with inclinations in the low, medium, and high inclination groupings, it was included in the sample of test systems for this work, until we reached ten systems total in each inclination grouping.
@@
The orbital parameters describing the shape (semimajor axis, eccentricity) and inclination of the planets’ orbits in each of the ten systems are listed in Table 2. We focus on the shape parameters and inclination here as these are the primary orbital parameters which will be used to initially characterize an exoplanet’s orbit. The semimajor axis is essential for constraining whether or not an observed exoplanet lies within the host star’s projected habitable zone, which is a defining requirement for deciding whether an observatory like HWO should allocate additional time and resources for follow-up observations (The LUVOIR Team, 2019).
Figure 3: Simulated detections for test case system ten across the low (a), medium (b), and high (c) inclination ranges (groups described in Section 2.3.1). Observations were spaced equally over one Earth year and each epoch of observation is represented by a different marker (circle, square, and triangle). Orbital parameters for the simulated systems are given in Table 2.
@@ 2.3.2 Simulating Detections
Examples of simulated detections are shown in Figure 3 for system ten. We assume observations are equally spaced to occur three times over a period of one Earth year. We choose to simulate three observations to find highly confused systems, and because previous work has demonstrated that the minimum number of observations required to sufficiently constrain a planet’s orbit to below 10% is three visits (Guimond & Cowan, 2019; Horning et al., 2019; Romero-Wolf et al., 2021). We assume each system is located at a distance of 10 parsecs from the observer with a Sun-like host star. The best potential target stars identified by Stapelfeldt & Mamajek (2023) (“Tier A”) for observation with HWO have distances that fall within a range of 3 to 19 parsecs. In our simulations, we also assume all planets are visible at all times for this initial study. This paints an overly optimistic picture of confusion rates and the ability to leverage photometry for high inclination systems; an investigation of null-detections is left for future work. Each simulated planet is prescribed a radius of 1 $R_{\earth}$ for the photometry calculations and a geometric albedo of 0.3. We prescribe these parameters to emulate observations of rocky Earth-like planets, which will be primary targets for HWO.

### v2

@@ 2 Approach
In this work, we investigate the influence of photometry on deconfusion by developing a photometry model and a new likelihood orbit-ranking scheme to augment the deconfuser. We demonstrate the utility of photometry for differentiating orbit assignments using a representative set of highly confused corner cases from three inclination groups (low, medium, and high inclination), assuming all planets are visible at all times. We adopt the definition of confusion from Ref. Pogorelyuk et al. [53], where a system is considered to be confused if the deconfuser provides two or more closely ranked orbit assignments per system. When fit by the deconfuser, orbit options may be considered closely ranked if they fall within the acceptable rms error tolerance, which is set to 0.05 AU in this work. After fitting orbits and ranking assignments based on the number of matched detections with the deconfuser, we use an additional ranking scheme to break ties between orbits if they have a more statistically likely brightness result. Figure 2 illustrates the combined flow of the algorithms presented in this work.
@@ 2.1 Photometry Model
The photometry ranking scheme begins with a photometry model at one bandpass, which returns the phase angle and brightness of a planet at any location on its orbit and calculates the brightness measured at the instrument for each detection. To determine the expected contrast of a planet detection, we first calculate the planet’s separation from its host star. The deconfuser provides the three-dimensional coordinates (x, y, z) of a simulated planet’s location (see Fig. 2, step 1). These coordinates are used to calculate the planet’s separation from its host star with the basic Euclidean distance formula. Then, we use the coordinates to determine the phase angle of the planet detection with respect to the observer (see Fig. 2, step 2). As described in Sec. 1, the phase angle is defined as the angle between the host star, exoplanet, and the observer. Given this geometry, the phase angle is calculated using Eq. (2),
@@
where $\vec{r}$ is the vector describing the planet’s separation from the host star and $\vec{s}$ is the vector describing the system’s separation from the observer. For the corner cases described here (details given in Table 2), the planetary system locations are set to a distance of ten parsecs, placing them well within the range of distances of the most accessible targets presented in NASA Exoplanet Exploration Program’s Mission Target Star List for HWO [42].
@@ 2.1.1 Phase Function
The phase angle calculated in Eq. (2) is used to determine the phase function of the planet at each orbital phase. For simplicity, we assume a Lambertian phase function and adopt the following equation (as presented in Ref. Brown [6]):
@@
Combining Eqs. (1) and (3) brings us to an “expected” brightness of an observation. This is the planet-star flux ratio that we would expect to receive at the observing system.
Assuming Lambertian reflectance is a simple approximation that is commonly chosen when modeling and comparing reflected light from planets (e.g., Refs. Zugger et al. [72], Fujii and Kawahara [17], Cowan et al. [11], Morgan et al. [47], De Cock et al. [12]), but the Lambertian assumption neglects the dependence of scattering on albedo, wavelength, and complex atmospheric properties (e.g., Refs. Greco and Burrows [21], Mayorga et al. [45]). For an initial study of how planetary phase variation can inform orbit differentiation, a Lambertian approximates the overall shape of a phase curve well to model phase angle variation. Future work will incorporate phase functions that describe more complex processes, such as the one- or two-term Henyey-Greenstein phase functions, which introduce additional parameters to model forward- and back-scattering processes [24, 28].
@@ 2.1.2 Noise Model
To consider more realistic observations and to obtain the likelihood of a planet detection, we developed a noise model to include detector parameters in the planet flux returned by the photometry model (see Fig. 2, step 3). We adopt the instrument parameters defined for the Roman Coronagraph Instrument, as Roman’s Coronagraph has been extensively modeled and tested in preparation for its expected launch in Fall 2026 [52]. The instrument parameters for the Roman Coronagraph Instrument are available online through IPAC at Caltech [55], and those we use in our code (from Phase C) are listed in Table 1.
@@ 2.1.3 Detector Model
We convert the planet flux density to the planet photon count rate at the telescope following the method described in Ref. Robinson et al. [54], which was developed with a telescope like Roman in mind and modeled on the work presented in Ref. Brown [6]. The detector noise function accepts the calculated planet photon rate for each observation and propagates it through a simplified Electron-Multiplying Charge-Coupled Device (EMCCD) detector model, which adds noise properties to the signal. The model includes shot noise, dark current, read noise, and EMCCD gain. The result is a planet signal with detector noise added for each simulated detection, which amounts to the output electron counts from the detector due to the incoming planet signal. For each orbit option returned by the deconfuser, we calculate both the “observed” noisy photometry and the “expected” noisy photometry of the simulated planet detection. The “observed” photometry simulates the instrument’s observation of a planet. In other words, this is the measured photometry that a user would provide to the algorithm for a real observation. The “expected” photometry is the value one expects to receive from a detected planet based on the assumed phase angle and planet parameters, including detector noise.
@@ 2.2 Photometry Likelihood Ranking
After we have included the detector noise for each observation, we implement a ranking scheme to sort confused orbit options based on the likelihood of the calculated photon rates per system option, given the assigned orbital parameters. We obtain the likelihood of each detection from the probability distribution of the brightness. This value is returned by Fig. 2, step 4.
As mentioned in Sec. 2.1, the detector noise model uses instrument parameters defined for Roman. To generate the probability distribution for each detections’ brightness, the detector noise model first draws samples from a Poisson distribution describing the expected number of photons reaching the detector. The incident photons are multiplied by the detector quantum efficiency before the dark current is added to the signal. We account for the EMCCD gain ($g$) by drawing samples from a Gamma distribution, following the model described in Ref. Hirsch et al. [25] and applied in LOWFSSim (the Roman Coronagraph Instrument optical model of low-order wavefront sensing and control [13]; https://github.com/nasa-jpl/lowfssim). The shape parameter of the gamma distribution is the number of input photoelectrons ($n_{ie}$), and the scale parameter ($\theta$) is
@@
The contributions to the noise model are a combination of three probability distributions: a Poisson distribution, $P(n_{p})$, describing the incident photons (Eq. 5), a Gamma distribution, $\Gamma(n_{ie},\theta)$, describing the EM register of the EMCCD (Eq. 6), and a normal distribution, $\mathcal{N}(R_{e^{\mathrm{-}}})$, describing the read noise (Eq. 7) to convert the signal to image values. The following equations describe these distributions.
@@
Where $q$ is the detector quantum efficiency, $\Delta t$ is the observation time, and $CIC$ is the detector clock-induced charge (see also Table 1). Combining Eqs. (5), (6), and (7) gives the probability of measuring a certain number of counts for a detection, $p(n_{c})$:
@@
The output of the noise model is the probability density function ($f(y|n_{p})$ in Eq. 10) describing the number of photoelectrons read out by the detector, from which we can determine the likelihood. The likelihood of a detection, given the orbital parameters of the orbit fit, is obtained by comparing where the observed brightness falls along this probability density function. We assume each planet detection is an independent event, so the likelihood of measuring all planet detections given the parameters of the matched orbits is the product of each individual detection. The resulting likelihood, $L_{\mathrm{system}}$, is used as a new ranking score for each set of confused orbit options per system. A more detailed mathematical description of this process is given in Appendix A. We note that in some cases, it may be beneficial to examine likelihoods on a per-planet basis to avoid a biased likelihood due to one poorly fit planet in the system. We expand on this in more detail in Section 4.
@@ 2.3 Simulated Systems
This section describes our method for identifying representative highly confused systems to demonstrate the photometry ranking scheme on. We simulate ten test case systems across three inclination groupings, low ($i<45^{\circ}$), medium ($45^{\circ}<i<70^{\circ}$), and high ($i>70^{\circ}$), to investigate how exoplanet photometry may support deconfusion of multi-planet systems across viewing inclinations. The ten simulated systems are each different systems, and only inclination is changing across the three inclination groupings, creating a total of thirty cases for study. Each of the simulated systems contains three planets. We chose the first ten highly confused cases from a random sample of simulated systems. The methods for simulating and identifying these confused systems are described in Sec. 2.3.1 below.
We restrict this analysis to three-planet systems to demonstrate the photometry ranking scheme based on Ref. Pogorelyuk et al. [53]’s findings regarding confusion in systems with three or more planets. The cases in this work were intentionally chosen because they are clearly challenging cases. That is, each of the simulated systems examined here results in confusion every time they are passed through the initial version of the deconfuser (without photometry), at each of the low, medium, and high inclinations. The purpose of this work is not to present large number statistics, but rather to explore if photometry is sufficient for deconfusing these representative known highly confused scenarios. We also demonstrate the photometry ranking scheme on a few Solar System planets in Appendix B.
The simulated systems presented here are not explicitly evaluated for dynamical stability. We simulate systems with orbital separations greater than 0.3 AU, with mutual inclinations and low eccentricities. For planets with masses similar to Earth, separations of 0.3 AU would place planets at separations well outside several times the Hill radii of the other planets in the system (see e.g., Ref. Gladman [20] or Dulz et al. [14] for discussion of using mutual Hill radii as a basic stability criterion). Dynamical simulations by Ref. Smith and Lissauer [59] indicate that Earth-mass planets in three-planet systems are stable for much longer than one billion years if their initial orbital separation is at least 7.5 times their mutual Hill radii. Simulations also show that stable multi-planet systems generally have lower median eccentricities and mutual inclinations [23].
@@ 2.3.1 Identifying Confused Systems
The simulated systems were chosen by first identifying confused systems with inclinations of less than 45. Due to the low probability of confusion for systems with $i<45^{\circ}$, it is more challenging to identify confused low inclination cases. For this reason, we identify test cases for this work starting with low inclination simulations and working our way up. Systems with inclinations less than $\sim$20 are also rarely confused [53]. Thus, there are no systems in our sample set which have an inclination less than 20.
Low inclination systems were simulated with three co-planar planets with random orbital parameters drawn from uniform distributions with semimajor axes between 0.5 and 6.0 AU, eccentricities less than 0.1, inclinations less than 45, and minimum orbital separations between planets of 0.3 AU. These systems were simulated using the system generation function (sample_planets.random_planet_elements) contained within the deconfuser framework.
Once a low inclination system was found to be confused with the first version of the deconfuser, that system was passed through the deconfuser again with only the inclination changed. Updated inclination values were randomly sampled from uniform arrays ranging between 45 to <<70 until a confused system was identified in the medium inclination range through empirical trial and error testing. The same method was used for the high inclination grouping, but inclinations were sampled between 70 and 90. If a system returned confused options with inclinations in the low, medium, and high inclination groupings, it was included in the sample of test systems for this work, until we reached ten systems total in each inclination grouping. By varying only inclination across otherwise identical systems, we are also able to assess whether the inclusion of brightness information more strongly supports deconfusion for low, medium, or high inclination systems.
@@
The orbital parameters describing the shape (semimajor axis, eccentricity) and inclination of the planets’ orbits in each of the ten systems are listed in Table 2. We focus on the shape parameters and inclination here as these are the primary orbital parameters that will be used to initially characterize an exoplanet’s orbit. The semimajor axis is essential for constraining whether or not an observed exoplanet lies within the host star’s projected habitable zone, which is a defining requirement for deciding whether an observatory like HWO should allocate additional time and resources for follow-up observations [64].
Figure 3: Simulated detections for test case system ten across the low (left), medium (center), and high (right) inclination ranges (groups described in Sec. 2.3.1). Observations were spaced equally over one Earth year and each epoch of observation is represented by a different marker (circle, square, and triangle). Orbital parameters for the test case systems are given in Table 2.
@@ 2.3.2 Simulating Detections
Examples of simulated detections are shown in Fig. 3 for system ten. We assume observations are equally spaced to occur three times over a period of one Earth year. We choose to simulate three observations to find highly confused systems, and because previous work has demonstrated that the minimum number of observations required to sufficiently constrain a planet’s orbit to below 10% is three visits [22, 26, 56]. We assume each system is located at a distance of 10 parsecs from the observer with a Sun-like host star. The best potential target stars identified by Ref. Mamajek and Stapelfeldt [42] (“Tier A”) for observation with HWO have distances that fall within a range of 3 to 19 parsecs. In our simulations, we also assume all planets are visible at all times for this initial study. This paints an overly optimistic picture of confusion rates and the ability to leverage photometry for high inclination systems; an investigation of null-detections is left for future work. Each simulated planet is prescribed a radius of 1 $R_{\oplus}$ for the photometry calculations and a geometric albedo of 0.3. We prescribe these parameters to emulate observations of rocky Earth-like planets, which will be primary targets for HWO.